OpenStudy (anonymous):

Find the center, vertices, and foci of the ellipse with equation 2x^2 + 7y^2 = 14.

OpenStudy (mertsj):

1. Divide both sides by 14 2. Put the equation into the form: \[\frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1\]

OpenStudy (mertsj):

Center is (h,k) Vertices are "a" units to the right and to the left of the center. \[a^2-b^2=c^2\]

OpenStudy (mertsj):

Find c. Foci are c units to the right and c units to the left of the center.

OpenStudy (anonymous):

after step one i got lost :(

OpenStudy (mertsj):

What did you get after you did Step 1?

OpenStudy (anonymous):

(2x^2)+(7y^2)/14 =1?

OpenStudy (mertsj):